OpenStudy (anonymous):

Find the slope of the line that passes through the points (3, -6) and (6, 12).

OpenStudy (alexandervonhumboldt2):

slope formula is (y1-y2)/(x1-x2)=(-6-12)/(3-6)=(-18)/(-3)=18/3=???

OpenStudy (alexandervonhumboldt2):

(x1, y1) and (x2, y2)are given coordinates

OpenStudy (anonymous):

Find the value of x so that the line passing through (-2, 7) and (x, 2) has a slope of

OpenStudy (anonymous):

-1/3

OpenStudy (alexandervonhumboldt2):

(7-2)/(-2-x)=-1/3 (-2-x)=-3*5 -x=-15+2 x=13

OpenStudy (anonymous):

Find the slope of a line that has an x-intercept of -2 and a y-intercept of 6.